OpenStudy (anonymous):

Unit tangent vector problem: Vector r(t) = for t is great than or equal to 1. Find unit tangent vector.

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

definition : \(\large \frac{r'(t)}{|r'(t)|}\)

ganeshie8 (ganeshie8):

to take the derivative of a vector function, simply take derivative of each component

OpenStudy (anonymous):

I think I am having trouble factoring the magnitude on bottom.

OpenStudy (anonymous):

I have T(t) = <1, 0, -2/t^2> / ( 1 + 4/t^4 )

ganeshie8 (ganeshie8):

looks good

OpenStudy (anonymous):

sorry T(t) = <1, 0, -2/t^2> / ( 1 + 4/t^4 )^1/2

ganeshie8 (ganeshie8):

oh yah

OpenStudy (anonymous):

my book shows <t^2, 0, -2> (t^4 + 4)^1/2 ?

ganeshie8 (ganeshie8):

thats right

ganeshie8 (ganeshie8):

\(\large \frac{<1, 0, -2/t^2>}{\sqrt{1 + 4/t^4}} \) \(\large \frac{<1, 0, -2/t^2>}{\sqrt{(t^4 + 4)/t^4}} \)

ganeshie8 (ganeshie8):

\(\large \frac{<1, 0, -2/t^2>}{\sqrt{(t^4 + 4)}/t^2} \)

ganeshie8 (ganeshie8):

multiply top and bottom wid \(t^2\)

ganeshie8 (ganeshie8):

\(\large \frac{t^2<1, 0, -2/t^2>}{\sqrt{(t^4 + 4)}} \)

OpenStudy (anonymous):

so did you first factor out t^4 ? O couldnt quite tell.

ganeshie8 (ganeshie8):

we have below in the denominator :- \(\large 1 + \frac{4}{t^2}\) right ?

ganeshie8 (ganeshie8):

to add fractions, u need to have common-denominator

ganeshie8 (ganeshie8):

\(\large 1 + \frac{4}{t^4}\) \(\large 1\times\frac{t^4}{t^4} + \frac{4}{t^4}\) \(\large \frac{t^4}{t^4} + \frac{4}{t^4}\)

OpenStudy (anonymous):

got it.

ganeshie8 (ganeshie8):

now that the denominators are same, we can add the numerators directly : \(\large 1 + \frac{4}{t^4}\) \(\large 1\times\frac{t^4}{t^4} + \frac{4}{t^4}\) \(\large \frac{t^4}{t^4} + \frac{4}{t^4}\) \(\large \frac{t^4+4}{t^4} \)

OpenStudy (anonymous):

I see whre your going with it (; Can I show you one more ting fro a different eq. you mght know how to simplify in a similiar way? 4sin^2(2x) + 16cos^2(2x)

ganeshie8 (ganeshie8):

thats easy, write 16 as 4 + 12

ganeshie8 (ganeshie8):

4sin^2(2x) + 16cos^2(2x) 4sin^2(2x) + 4cos^2(2x) + 12cos^2(2x) 4(sin^2(2x) + cos^2(2x) ) + 12cos^2(2x) 4(1) + 12cos^2(2x) 4 + 12cos^2(2x)

ganeshie8 (ganeshie8):

i dont think it simplifies anymore

OpenStudy (anonymous):

Thats it and thank you. What time is it in India?

ganeshie8 (ganeshie8):

np :) 1:04 pm .... going for lunch.... cya must be very late over there. lol u should goto sleep :)

OpenStudy (anonymous):

Its 11 36 pm (; No class tomorrow. Have a good day.